OpenStudy (anonymous):

find the solution to the equation 64^(x - 3) = 4^2x

OpenStudy (lgbasallote):

\[\large \Rightarrow 64^{(x-3)} = 4^{2x}\] \[\large \Rightarrow 4^{3(x-3)} = 4^{2x}\] \[\large \Rightarrow 3(x-3) = 2x\] \[\large \Rightarrow 3x - 9 = 2x\] do you get the process??/

OpenStudy (anonymous):

No not really. Like how did you get "4^3(x - 3)" ?

OpenStudy (lgbasallote):

\[\Large 64 = 4 \times 4 \times 4 = 4^3\]

OpenStudy (lgbasallote):

so i subbed \(4^3\) into 64..got it now?

OpenStudy (anonymous):

Okay, What is this method called?

OpenStudy (lgbasallote):

which method? the whole one? or turning 64 into 4^3?

OpenStudy (anonymous):

Yeah, I have to simplify the equation till I get x = -9?

OpenStudy (anonymous):

The whole method for the equation.

OpenStudy (lgbasallote):

careful... 3x - 9 = 2x 3x - 2x = 9

OpenStudy (lgbasallote):

not negative

OpenStudy (anonymous):

okay. so it's x = 9 ?

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

idk the formal name of the method..but it has something to do with putting the bases the same

OpenStudy (anonymous):

Those are called exponent rules. If: \[5^{6x}=5^{18x}\] Then 6x=18x

OpenStudy (anonymous):

exponential rules rather

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem, just google them to get familiar with other rules.

OpenStudy (theviper):

I m with lgbasallote!:)